Fatty is three years younger than Annie. Annie is twice as old as Mitch. The sum of their ages is 42. How old each?
Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52781)

.

Mitch + Annie + Fatty  = 42


  M   +  2M   + (2M-3) = 42


       5M              = 42 + 3 = 45

        M                       = 45/5 = 9.


ANSWER.  9 years (Mitch);  18 years (Annie),  and  18-3 = 15 (Fatty)

Solved.



Answer by josgarithmetic(39617)

PERSON              AGE
Fatty                2M-3
Annie                2M
Mitch                M
SUM                 42

.
.